OpenStudy (anonymous):

Solve for x: 1/4 (2x − 14) = 4

OpenStudy (anonymous):

hello! welcome to open study =)

OpenStudy (anonymous):

ok so to find "x" what is our objective?

OpenStudy (anonymous):

@ spssss

OpenStudy (anonymous):

correct =) so, our objective is to isolate x...

OpenStudy (anonymous):

ok, I see you closed the question. When you close a question that means that it is "finished". But, do you still need help?

OpenStudy (anonymous):

yes my lesson is to find the value of x with fraction in the problem ill give you the options @ScienceQs

OpenStudy (anonymous):

yes i do

OpenStudy (anonymous):

i do need help i havenever used this before

OpenStudy (anonymous):

@ScienceQs

OpenStudy (anonymous):

sure =). Our first step is to get x free of the parentheses. How should we do that?

OpenStudy (anonymous):

im not sure in the lessson videos and practice there wasnt a question like this its diffrent and my guess is to add 14 to the -14 in the parente.

OpenStudy (anonymous):

@ScienceQs

OpenStudy (anonymous):

ok. if you could give me a copy of the lesson on .doc or pdf that would be great. Our first step is to multiply 1/4 by everything that is in the parenthesis

OpenStudy (anonymous):

its on florida virtual school idk if i can @ScienceQs

OpenStudy (anonymous):

and i wasnt sure if i was suppoesed to multiply what was in parentese or not thats why i couldnt solve it. @ScienceQs

OpenStudy (anonymous):

ah, ok. parenthesis in math is just like the * in math. It means to multiply. For example, 2(3) = 5, and 2*3=5.

OpenStudy (anonymous):

how is that i thought 2*3=6

OpenStudy (anonymous):

@ScienceQs

OpenStudy (anonymous):

oops sorry i meant 6...i was typing fast. don't worry, I'll be more careful now! I was talking with somebody at the same time

OpenStudy (anonymous):

ok let's see: 1/4(2x-14) = 4 1/2x - 3.5 = 4 What to do next?

OpenStudy (anonymous):

@spssss I know, but I'm asking, what do you think should we do next?

OpenStudy (anonymous):

do we muliply 1/2?@ScienceQs

OpenStudy (anonymous):

@ScienceQs

OpenStudy (anonymous):

nope. 1/2x is "one-half-of-x". good try, though. What we need to do is get other numbers out of the way. To do that, we need to get rid of that -3.5, by adding it to both sides. Like this: 1/2x -3.5 = 4 +3.5 to both sides 1/2x = 7.5

OpenStudy (anonymous):

Now, we want a whole x by itself. What do you think we should do?

OpenStudy (anonymous):

i types that and thought it was worng :c

OpenStudy (anonymous):

well now we divide 1/2 dont we. @ScienceQs

OpenStudy (anonymous):

i mean multiply

OpenStudy (anonymous):

no i mean combine like term to make it x=9 @ScienceQs

OpenStudy (anonymous):

Nice idea, but we can't combine like terms, because 7.5 and x are not like terms. If it was 4x and x, then you would be right! But it's not that easy for us. What we need to do is multiply both sides by 2, like so: 1/2x = 7.5 *2 to both sides 1x = 15 1x is rewritten as x x = 15

OpenStudy (anonymous):

IM SO DUMB thank you so much now i have a reference ill copy some of this down in my notes.!!!!!!!

OpenStudy (anonymous):

Can you help me with another quesion?

OpenStudy (anonymous):

no, you are not dumb! You are very bright. This is new to you, so of course you wouldn't get it first time. I enjoy math, but it does get very tricky sometimes. I get confused very often when I'm learning something new: like right now, I'm learning matrices, and I have no idea how to solve them! That's why I'm so thankful for the nice people on Open Study, and for my teacher/tutor.

OpenStudy (anonymous):

Allright, but before we move on to another question, let's check the one we just did to make sure that it is correct and no mistakes =)

OpenStudy (anonymous):

Solve for x: 2/3 x + 1 = 27 and sure

OpenStudy (anonymous):

@ScienceQs i have 2 questons and then i will be done with this lesson i have lots of trouble.

OpenStudy (anonymous):

k: 1/4 (2x − 14) = 4 we know (or think we know) that x =15. solve it and I'll check your work =) then let's move on to the next Q

OpenStudy (anonymous):

i have 1 left i will give you the equation later.

OpenStudy (anonymous):

here 1/2x-3.5=4 1/2x=7.5 x=15 the most diffficuly part is finding out what to do when. @ScienceQs

OpenStudy (anonymous):

it cut off half my work :c

OpenStudy (anonymous):

sure. have you gone over PEMDAS yet?

OpenStudy (anonymous):

yes but it get confusing with fraction @ScienceQs

OpenStudy (anonymous):

the eqaution ive tryed 100 times is 5/2x+15/2=15 what i keep getting is 10x=30 = 30 and i know it incorrect @ScienceQs

OpenStudy (anonymous):

ok. let me just give you a little tip: I use a different thing than PEMDAS. I don't know if it would help you better or not, but it did for me so maybe it will for you, too. It's called G-PAW and it goes like this: G- grouping symbols. see those grouping symbols like shields, and you have to destroy them to get at the numbers inside! It doesn't matter what kind of shields you're up against, they need to go down! P- powers. solve them if they're there. A-and...multiply, divide, add, subtract (in order, as necessary) W- work left to right

OpenStudy (anonymous):

allright, now for your equation...is there a way i can see it on paper? the dividing symbols confuse me, like i don't know how much we're dividing here. maybe you could write it out like it is, scan it, and put on your pc as jpg file then attatch the file?

OpenStudy (anonymous):

its not divideing there fraction plus printer / scanner is broken

OpenStudy (anonymous):

Ok, for 5/2x+15/2=15 let's multiply 1/5 to both sides. you know what this does, right?

OpenStudy (anonymous):

i reset exam and dont hAVE THIS equation anymore thanks for the help

OpenStudy (anonymous):

would you still like to work through the problem together? It would be for good practice.

OpenStudy (anonymous):

you are welcome!